# JL1-CD: A New Benchmark for Remote Sensing Change Detection and a Robust Multi-Teacher Knowledge Distillation Framework

### v2

@@ IV-C Loss Function
FC-EF[10] CNN 1.353 12.976 1e-3 - LinearLR 8 3090
FC-Siam-Conc[10] CNN 1.548 19.956 1e-3 - LinearLR 8 3090
FC-Siam-Diff[10] CNN 1.352 17.540 1e-3 - LinearLR 8 3090
STANet-Base[22] ResNet-18 12.764 70.311 1e-3 5e-3 LinearLR 8 3090
@@
FarSeg (ResNet-18)[48] 16.965 76.845 1e-3 1e-3 LinearLR 16 3090
ChangeStar[24] UPerNet (ResNet-18)[49] 13.952 55.634 1e-3 1e-4 LinearLR 8 3090
MiT-b0 3.847 11.380 6e-5 1e-3 LinearLR 8 3090
@@
TABLE III: Experimental Results on JL1-CD Test Set
Method Strategy mIoU mAcc mPrecision mFscore Method Strategy mIoU mAcc mPrecision mFscore
STANet (Base) - 66.76 81.71 74.73 74.73 IFN - 71.25 78.91 84.53 77.33
O-P 64.56 78.47 78.47 71.25 O-P 71.06 78.37 84.28 77.21
MTKD 67.92 82.07 76.24 75.10 MTKD 72.72 80.28 84.66 78.80
SNUNet (c16) - 68.97 74.87 85.06 75.25 BIT - 67.22 74.47 83.71 73.37
O-P 71.39 78.60 83.36 77.98 O-P 69.41 76.29 84.02 75.77
MTKD 71.12 78.27 84.96 77.56 MTKD 68.86 75.49 84.71 74.88
ChangeStar (FarSeg) - 69.47 75.58 84.46 75.57 ChangeStar (UPerNet) - 64.85 69.18 88.26 70.19
O-P 68.87 74.74 84.90 74.86 O-P 64.68 69.05 87.23 70.08
MTKD 69.14 76.49 82.09 75.41 MTKD 65.10 70.26 87.69 70.58
ChangeFormer (MiT-b0) - 73.51 80.46 86.33 79.70 ChangeFormer (MiT-b1) - 73.05 79.70 86.95 79.22
O-P 72.58 79.16 86.33 78.79 O-P 73.45 79.19 87.45 79.41
MTKD 73.25 79.20 87.15 79.30 MTKD 73.92 80.43 86.89 80.18
TinyCD - 71.04 78.77 83.05 77.74 HANet - 63.64 69.77 83.43 69.39
O-P 72.22 79.93 83.49 78.76 O-P 69.05 76.53 83.05 75.66
MTKD 72.55 80.98 83.17 79.26 MTKD 67.67 74.39 84.38 73.92
Changer (MiT-b0) - 74.85 81.84 86.09 80.98 Changer (MiT-b1) - 75.94 81.99 87.74 81.93
O-P 75.29 81.40 87.06 81.32 O-P 75.42 81.67 87.13 81.43
MTKD 75.35 81.76 87.18 81.28 MTKD 76.15 82.85 86.98 82.13
Changer (r18) - 68.37 75.15 83.43 74.54 Changer (s50) - 62.31 69.23 80.91 67.83
O-P 70.76 77.42 83.86 77.01 O-P 71.80 79.76 83.15 78.23
MTKD 69.45 77.26 81.50 75.86 MTKD 62.96 69.65 81.76 68.52
LightCDNet (s) - 66.70 73.21 83.45 72.46 CGNet - 73.37 80.31 85.33 79.65
O-P 70.19 77.43 83.99 76.16 O-P 72.95 79.71 85.50 79.12
MTKD 65.99 72.44 83.86 71.48 MTKD 73.82 80.32 86.33 79.91
BAN (ViT-L) - 73.54 79.54 87.89 79.47 TTP - 75.05 80.24 89.82 80.76
O-P 73.61 79.17 88.10 79.45 O-P 76.69 83.48 87.27 82.52
MTKD 73.95 80.26 87.12 79.92 MTKD 76.85 82.99 88.05 82.56
BAN (ViT-B) - 73.30 80.36 85.91 79.47 BAN (ViT-B-IN21K) - 74.69 81.09 87.14 80.75
O-P 72.47 78.78 86.31 78.58 O-P 73.50 79.98 86.25 79.50
FC-EF - 57.08 61.90 86.40 61.28 FC-Siam-Conc - 63.79 69.54 84.77 69.19
O-P 49.59 53.30 95.54 51.47 O-P 60.25 63.84 91.19 64.72
FC-Siam-Diff - 61.30 66.03 86.45 66.34
O-P 56.49 60.05 91.64 60.57
V EXPERIMENT
V-A Dataset Description
We first conduct experiments on our JL1-CD dataset. Additionally, to validate the robustness of the proposed O-P strategy and MTKD framework, we further perform experiments on the SYSU-CD dataset [47]. The detailed information for these two datasets is as follows:
1) JL1-CD Dataset: As described in Section III, the JL1-CD dataset consists of 5,000 pairs of high-resolution images, with a resolution of 0.5-0.75 meters and image size of 512 × 512. In the competition, the first 4,000 image pairs are used for training, and the remaining 1,000 image pairs are used for testing (with the ground truth labels being unavailable during the competition). To ensure sufficient data for training, we use the first 100 image pairs as the validation set, the next 3,900 image pairs as the training set, and the remaining 1,000 image pairs for testing. As shown in Fig. 5, our data split is reasonable, as the CAR distributions of the three sets are very similar.
2) SYSU-CD Dataset: As illustrated in Fig. 6, the SYSU-CD dataset is also a challenging dataset with a very large CAR range, so we choose this dataset as the second benchmark to validate the robustness of our proposed methods. The SYSU-CD dataset contains 12,000 pairs for training, 4,000 pairs for validation, and 4,000 pairs for testing, with each image having a resolution of 0.5 meters and a size of 256 × 256.
Figure 7: Visual comparison on the JL1-CD dataset. Each row, from top to bottom, represents the following: image at time 1, image at time 2, ground truth, output from the original model, output from the O-P strategy, and output from the MTKD framework. Red denotes missed detections (FN), while blue indicates false alarms (FP). The selected algorithms are: (a) BAN-ViT-L, (b) BIT, (c) TTP, (d) SNUNet, (e) IFN, (f) Changer-MiT-b1, (g) ChangeFormer-MiT-b1, (h) TinyCD, and (i) CGNet.
V-B Benchmark Methods
To comprehensively verify the validity of our JL1-CD dataset and the superiority of the proposed O-P strategy and MTKD framework, we conduct extensive experiments on existing benchmark algorithms. The selected models, along with their corresponding backbones, parameter sizes, and computational complexities, are summarized in Table II. Based on the backbone architecture, the models are categorized into three groups: Alice Blue for CNN-based models, Light Cyan for Transformer-based models, and Lavender Blue for FM-based models, encompassing almost all mainstream architectures. FLOPs is calculated with an input image of size 512 × 512. As shown in the table, the selected models span a wide range of sizes, from lightweight models with fewer than 1M parameters, such as TinyCD and LightCDNet, to the latest SOTA model TTP with over 360M parameters. The inclusion of these lightweight models is crucial, as improving their detection performance without increasing inference resource consumption is of significant importance for deployment in edge devices and scenarios requiring real-time processing. This wide range of model sizes also allows us to verify the universality of the O-P and MTKD methods across models with different backbones and scales.
V-C Evaluation Metrics
The common evaluation metrics for CD models include Intersection over Union (IoU), accuracy, precision, recall, and F1-score. IoU measures the overlap between the detected change region and the ground truth. Accuracy reflects the overall correctness of the model. Precision indicates the false positive rate of the model, recall reflects the false negative rate, and F1-score balances both precision and recall. A higher F1-score indicates better detection performance.
However, given the large CAR range in the JL1-CD dataset, both change and non-change regions are equally important. Therefore, we choose the averaged versions of the aforementioned metrics, which are calculated as follows:
mIoU $=\frac{1}{2}\left(\text{IoU}_{0}+\text{IoU}_{1}\right)$ (8)
$=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FP}+\text{FN}}+\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}}\right)$
mPrecision $=\frac{1}{2}\left(\text{Precision}_{0}+\text{Precision}_{1}\right)$
$=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FN}}+\frac{\text{TP}}{\text{TP}+\text{FP}}\right)$
mAcc $=\frac{1}{2}\left(\text{Recall}_{0}+\text{Recall}_{1}\right)$
$=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FP}}+\frac{\text{TP}}{\text{TP}+\text{FN}}\right)$
mFscore $=\frac{1}{2}\left(\text{Fscore}_{0}+\text{Fscore}_{1}\right)$
$=\frac{1}{2}\sum_{i=0}^{1}\frac{2\times\text{Precision}_{i}\times\text{Recall}_{i}}{\text{Precision}_{i}+\text{Recall}_{i}}$
where TP, FP, TN, and FN represent true positives, false positives, true negatives, and false negatives, respectively. The averaged accuracy and recall are equivalent, so we only use mAcc for consistency in subsequent experiments.
It is important to note that in the MMSegmentation[50] toolbox, these metrics are computed based on the total number of pixels across all predicted label images. In this paper, however, to align with the competition requirements, we first calculate these metrics for each individual image and then compute the average of the results across all images to obtain the final outcome.
Figure 8: mIoU of HANet, ChangeFormer-MiT-b1, and TTP across different CAR ranges. The first and second rows show results on the validation and test sets, respectively. In each plot, the left y-axis represents CAR size, and the right y-axis represents mIoU.
V-D Implementation Details
@@
All algorithms are trained and tested using the PyTorch-based OpenCD Toolbox [51]. To ensure fair comparisons and to clearly assess the contributions of the Origin-Partition (O-P) strategy and the Multi-Teacher Knowledge Distillation (MTKD) framework to the performance improvement of the CD models, we adopt consistent parameter settings and hardware conditions for all models ($\mathcal{M}_{O}$, $\mathcal{M}_{T_{L}}$, $\mathcal{M}_{T_{M}}$, $\mathcal{M}_{T_{S}}$, and $\mathcal{M}_{S}$) across the various algorithms, as described below:
@@
    Training begins with a warm-up phase of 1k iterations (5 epochs for TTP), during which the learning rate (LR) is linearly increased from 1e-6 to the initial LR value, as specified in Table II. Afterward, the LR is linearly decayed to 0 as training progresses (TTP employs a CosineAnnealing decay schedule).
@@
    The HANet, CGNet, BAN (ViT-L), and TTP models are trained on the NVIDIA A800 server, while other models are trained on the NVIDIA RTX 3090 server. All models are tested on the A800 server.
@@
When partitioning the dataset, we set the thresholds $th_{1}=0.05$ and $th_{2}=0.2$ by empirical, which ensures a balanced distribution of samples across the partitions. The model is saved every 1k iterations (5 epochs for TTP), and the checkpoint with the highest mIoU value on the validation set is selected for testing.
@@
For the training of $\mathcal{M}_{T_{S}}$, due to the varying distributions of the change maps and the different magnitudes of $L_{CE}$ across modelss, a grid search is performed over the set $\{1e-5,5e-5,1e-4,5e-4,1e-3,5e-3,1e-2\}$ to determine the optimal distillation loss weight $\lambda$ for each model. More configuration details are summarized in Table II.
TABLE IV: Comparison of Detection Results on Change and No-Change Classes
Method Class IoU Acc Precision Fscore
unchanged +0.24 +0.29 +0.04 +0.12
IFN changed +2.71 +2.44 +0.22 +2.82
unchanged +0.10 -0.60 +0.65 +0.06
SNUNet (c16) changed +4.21 +7.38 -0.86 +4.54
unchanged +0.21 -0.02 +0.24 +0.18
ChangeFormer (MiT-b0) changed -0.74 -2.51 +1.41 -0.99
unchanged +0.07 +0.12 -0.01 +0.05
ChangeFormer (MiT-b1) changed +1.68 +1.35 -0.11 +1.86
unchanged +0.30 +0.29 +0.08 +0.20
TinyCD changed +2.72 +4.13 +0.16 +2.85
unchanged +0.21 +0.32 -0.14 +0.19
Changer (MiT-b0) changed +0.80 -0.48 +2.32 +0.41
unchanged +0.02 +0.09 -0.04 -0.01
Changer (MiT-b1) changed +0.41 +1.63 -1.47 +0.42
unchanged +0.02 -0.03 -0.04 -0.06
CGNet changed +0.88 +0.03 +2.04 +0.59
unchanged +0.12 +0.23 -0.09 +0.07
BAN (ViT-L) changed +0.70 +1.21 -1.46 +0.82

### v3

@@ Appendix B Benchmark Methods
FC-EF[2] CNN 1.353 12.976 1e-4 1e-4 LinearLR 8 3090
FC-Siam-Conc[2] CNN 1.548 19.956 1e-4 1e-4 LinearLR 8 3090
FC-Siam-Diff[2] CNN 1.352 17.540 1e-4 1e-4 LinearLR 8 3090
STANet[4] ResNet-18 12.764 70.311 1e-3 5e-3 LinearLR 8 3090
@@
ChangeStar-FarSeg[6] ResNet-18 16.965 76.845 1e-3 1e-4 LinearLR 16 3090
ChangeStar-UPerNet[6] ResNet-18 13.952 55.634 1e-3 1e-4 LinearLR 8 3090
MiT-b0 3.847 11.380 6e-5 1e-4 LinearLR 8 3090
@@
Appendix C Implementation Details
@@
All models are trained and evaluated with the OpenCD Toolbox built on PyTorch [38]. To ensure fair comparisons and to clearly assess the contributions of the Origin-Partition (O-P) strategy and the Multi-Teacher Knowledge Distillation (MTKD) framework to the performance improvement of the CD models, we adopt consistent parameter settings and hardware conditions for all models ($\mathcal{M}_{O}$, $\mathcal{M}_{T_{L}}$, $\mathcal{M}_{T_{M}}$, $\mathcal{M}_{T_{S}}$, and $\mathcal{M}_{S}$) across the various algorithms, as described below:
@@
    Training begins with a warm-up phase of 1k iterations (5 epochs for TTP), during which the learning rate (LR) is linearly increased from 1e-6 to the initial LR value, as specified in Table 6. Afterward, the LR is linearly decayed to 0 as training progresses (TTP employs a CosineAnnealing decay schedule).
@@
    The HANet, CGNet, BAN, and TTP models are trained on the NVIDIA A800 server, while other models are trained on the NVIDIA RTX 3090 server. All models are tested on the A800 server.
@@
When dividing the training set into three partitions, we empirically set the thresholds as $th_{1}=0.05$ and $th_{2}=0.2$. For the case using only two teacher models, the threshold is set to $0.10$. The model is saved every 1k iterations (5 epochs for TTP), and the checkpoint with the highest mIoU value on the validation set is selected for testing.
@@
For the training of $\mathcal{M}_{T_{S}}$, due to the varying distributions of the change maps and the different magnitudes of $L_{CE}$ across models, a grid search is performed to determine the optimal distillation loss weight $\lambda$ for each model. More configuration details are summarized in Table 6.
Appendix D Evaluation Metrics
The common evaluation metrics for CD models include Intersection over Union (IoU), Precision (Prec), Recall (Rec), and F1-score (Fscore). IoU measures the overlap between the detected change region and the ground truth. Accuracy reflects the overall correctness of the model. Precision indicates the false positive rate of the model, Recall reflects the false negative rate, and F1-score balances both of them. A higher F1-score indicates better detection performance. Given the large CAR range in the JL1-CD dataset, both change and non-change regions are equally important. Therefore, we choose the averaged versions of the aforementioned metrics, which are calculated as follows:
$\text{mIoU}=\frac{1}{2}\left(\text{IoU}_{0}+\text{IoU}_{1}\right)=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FP}+\text{FN}}+\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}}\right),$ (6)
$\text{mPrec}=\frac{1}{2}\left(\text{Prec}_{0}+\text{Prec}_{1}\right)=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FN}}+\frac{\text{TP}}{\text{TP}+\text{FP}}\right),$
$\text{mRec}=\frac{1}{2}\left(\text{Rec}_{0}+\text{Rec}_{1}\right)=\frac{1}{2}\left(\frac{\text{TN}}{\text{TN}+\text{FP}}+\frac{\text{TP}}{\text{TP}+\text{FN}}\right),$
$\text{mFscore}=\frac{1}{2}\left(\text{Fscore}_{0}+\text{Fscore}_{1}\right)=\frac{1}{2}\sum_{i=0}^{1}\frac{2\times\text{Prec}_{i}\times\text{Rec}_{i}}{\text{Prec}_{i}+\text{Rec}_{i}},$
where TP, FP, TN, and FN represent true positives, false positives, true negatives, and false negatives, respectively. To align with the “Jilin-1” Cup competition requirements, we first calculate these metrics for each individual image and then compute the average of the results across all images to obtain the final outcome.
Appendix E Additional Experiments on the JL1-CD Dataset
We summarize the quantitative results of 21 models on both change and no-change regions in Table 7 to provide a comprehensive evaluation of the JL1-CD dataset as a new benchmark and to further validate the effectiveness and generality of the MTKD framework. All models exhibit improvements in IoU and Fscore for change regions after applying MTKD.
Interestingly, we observe that MTKD tends to improve recall more than precision. For instance, the recall of the change region in the MTKD-enhanced TTP model improves by 5.69%, while precision drops by 3.99%. Nevertheless, the final F1-score still increases by 3.39%. This suggests that MTKD effectively reduces the model’s miss rate, and although it may increase false alarms for some algorithms, the overall performance gains are evident.
In Figure 8, we further visualize the mIoU performance of all benchmark methods across different CAR partitions. Consistent with the findings in Section 5, MTKD yields more notable improvements in partitions with low CAR values, while its impact on high-CAR partitions is generally less pronounced. This is exactly what we need: original models already perform well on large-scale changes (with mIoU often exceeding 70% in the highest CAR partition), but struggle with small-scale changes (where mIoU frequently falls below 20% in the lowest CAR partition). MTKD makes substantial contributions to improving the detection of small-scale changes.
Figure 9 presents additional visualization results, where it can be observed that missed detections (particularly for small object changes) occur most frequently across the original algorithms. This observation aligns with the quantitative metrics reported in Table 7. Furthermore, certain algorithms (e.g., Changer-ResNet-18, CGNet, and TTP) exhibit false alarms. Both the O-P and MTKD frameworks can effectively mitigate these issues.
Whether in Table 7 or Figure 8, we also observe that some models perform worse with the O-P strategy even compared to their original versions. As previously discussed, this can be attributed to partitioning errors caused by wrong CAR estimation, which may lead to suboptimal model selection during the second-stage inference. MTKD not only improves overall test efficiency compared to O-P but also enhances model robustness. In the following section, we provide a deeper analysis of failure scenarios under the O-P strategy.
Table 7: Experimental results on JL1-CD test set. To better illustrate the impact of the O-P strategy and MTKD framework on the detection performance of change and no-change classes, we report only the relative improvements in evaluation metrics compared to the original baseline model.
Method Strategy Class IoU Rec Prec Fscore Method Strategy Class IoU Rec Prec Fscore
unchanged 92.85 98.45 94.01 95.56 unchanged 93.95 97.89 95.70 96.30
- changed 21.30 25.36 78.79 27.00 - changed 33.64 41.19 73.84 42.08
unchanged -0.10 +0.48 -0.45 +0.08 unchanged +0.05 +0.38 -0.25 +0.13
O-P changed -13.95 -17.02 +18.20 -18.64 O-P changed -5.78 -10.02 +12.77 -7.59
unchanged +0.47 -0.95 +1.40 +0.39 unchanged -0.03 -0.42 +0.34 +0.06
FC-EF MTKD changed +9.12 +14.04 -10.40 +11.72 FC-EF MTKD changed +1.05 +2.49 -2.36 +1.32
unchanged 93.74 98.71 94.77 96.17 unchanged 92.27 94.22 97.64 95.41
- changed 28.86 33.34 78.13 36.51 - changed 41.26 69.21 51.82 52.63
unchanged +0.26 -0.44 +0.68 +0.26 unchanged -0.06 +0.79 -0.86 -0.04
O-P changed -1.00 -2.17 +8.48 -2.02 O-P changed -4.36 -7.28 -4.35 -5.49
unchanged +0.15 -0.28 +0.36 +0.10 unchanged +0.65 +0.74 -0.12 +0.33
FC-Siam-Diff MTKD changed +4.76 +7.36 -5.35 +6.21 STANet MTKD changed +1.66 -0.03 +3.13 +1.83
unchanged 94.59 97.34 96.93 96.76 unchanged 94.59 98.12 96.21 96.76
- changed 47.90 60.48 72.14 57.89 - changed 43.34 51.63 73.92 53.75
unchanged -0.17 -0.23 +0.12 -0.06 unchanged -0.06 -0.18 +0.03 -0.03
O-P changed -0.19 -0.85 -0.63 -0.17 O-P changed +4.91 +7.63 -3.44 +5.49
unchanged +0.24 +0.29 +0.04 +0.12 unchanged +0.10 -0.60 +0.65 +0.06
IFN MTKD changed +2.71 +2.44 +0.22 +2.82 SNUNet MTKD changed +4.21 +7.38 -0.86 +4.54
unchanged 94.14 97.34 96.47 96.50 unchanged 94.91 98.04 96.62 97.00
- changed 40.31 51.61 70.95 50.25 - changed 44.04 53.12 72.31 54.15
unchanged +0.13 +0.03 +0.14 +0.05 unchanged -0.48 -0.62 +0.02 -0.35
O-P changed +4.23 +3.59 +0.49 +4.75 O-P changed -0.73 -1.05 +0.86 -1.08
unchanged +0.42 +0.66 -0.10 +0.22 unchanged +0.13 -0.06 +0.19 +0.09
BIT MTKD changed +2.86 +1.37 +2.09 +2.79 ChangeStar (FarSeg) MTKD changed +0.65 +0.51 +2.44 +0.82
unchanged 94.36 98.44 95.67 96.57 unchanged 95.28 98.26 96.83 97.19
- changed 35.35 39.93 80.84 43.81 - changed 51.74 62.66 75.82 62.21
unchanged +0.15 -0.10 +0.20 +0.21 unchanged +0.13 -0.14 +0.24 +0.13
O-P changed -0.49 -0.16 -2.26 -0.42 O-P changed -1.99 -2.46 -0.22 -1.95
unchanged -0.08 -0.16 +0.08 -0.03 unchanged +0.26 -0.18 +0.47 +0.15
ChangeStar (UPerNet) MTKD changed +0.57 +2.31 -1.20 +0.81 ChangeFormer (MiT-b0) MTKD changed +0.28 -0.60 +0.69 +0.40
unchanged 95.40 98.08 97.12 97.32 unchanged 94.23 97.10 96.54 96.43
- changed 50.70 61.31 76.78 61.13 - changed 47.85 60.44 69.55 59.05
unchanged +0.04 +0.17 -0.11 +0.00 unchanged +0.30 +0.39 +0.19 +0.29
O-P changed +0.77 -1.17 +1.12 +0.38 O-P changed +2.06 +1.92 +0.69 +1.76
unchanged +0.07 +0.12 -0.01 +0.05 unchanged +0.30 +0.29 +0.08 +0.20
ChangeFormer (MiT-b1) MTKD changed +1.68 +1.35 -0.11 +1.86 TinyCD MTKD changed +2.72 +4.13 +0.16 +2.85
unchanged 93.53 97.48 95.51 96.04 unchanged 95.37 97.55 97.57 97.20
- changed 33.76 42.06 71.35 42.74 - changed 54.32 66.13 74.61 64.76
unchanged +0.67 +0.21 +0.60 +0.51 unchanged +0.04 +0.31 -0.33 +0.10
O-P changed +10.14 +13.32 -1.37 +12.02 O-P changed +0.84 -1.19 +2.27 +0.59
unchanged +0.78 +0.28 +0.61 +0.53 unchanged +0.21 +0.32 -0.14 +0.19
HANet MTKD changed +7.27 +8.96 +1.29 +8.53 Changer (MiT-b0) MTKD changed +0.80 -0.48 +2.32 +0.41
unchanged 95.87 98.17 97.51 97.62 unchanged 94.07 97.58 96.13 96.39
- changed 56.01 65.81 77.96 66.24 - changed 42.66 52.71 70.73 52.68
unchanged -0.44 -0.10 -0.35 -0.30 unchanged +0.58 -0.16 +0.65 +0.38
O-P changed -0.61 -0.54 -0.86 -0.69 O-P changed +4.21 +4.71 +0.21 +4.56
unchanged +0.02 +0.09 -0.04 -0.01 unchanged +0.11 -0.33 +0.40 +0.07
Changer (MiT-b1) MTKD changed +0.41 +1.63 -1.47 +0.42 Changer (ResNet-18) MTKD changed +2.07 +4.57 -4.26 +2.57
unchanged 92.90 96.42 95.90 95.66 unchanged 93.52 96.65 96.52 96.05
- changed 31.72 42.03 65.92 40.00 - changed 35.63 45.66 69.71 44.21
unchanged +1.24 +0.68 +0.73 +0.73 unchanged +0.54 +0.31 +0.23 +0.33
O-P changed +17.73 +20.40 +3.75 +20.07 O-P changed +10.69 +12.24 +1.51 +11.73
unchanged +0.40 +0.48 -0.06 +0.27 unchanged +0.65 +0.87 -0.32 +0.45
Changer (ResNeSt-50) MTKD changed +0.89 +0.36 +1.77 +1.12 LightCDNet MTKD changed +2.19 +1.70 +1.81 +2.25
unchanged 95.34 98.23 96.93 97.30 unchanged 95.57 98.14 97.26 97.40
- changed 51.40 62.40 73.74 62.00 - changed 51.51 60.94 78.52 61.55
unchanged -0.38 -0.69 +0.24 -0.27 unchanged +0.01 +0.39 -0.46 +0.00
O-P changed -0.47 -0.53 +0.08 -0.78 O-P changed +0.14 -1.12 +0.87 -0.04
unchanged +0.02 -0.03 -0.04 -0.06 unchanged +0.12 +0.23 -0.09 +0.07
CGNet MTKD changed +0.88 +0.03 +2.04 +0.59 BAN MTKD changed +0.70 +1.21 -1.46 +0.82
unchanged 95.68 98.46 97.03 97.41
- changed 54.43 62.02 82.62 64.12
unchanged -0.14 -0.56 +0.39 -0.06
O-P changed +3.42 +7.03 -5.50 +3.58